OpenStudy (anonymous):

Use the quadratic formula to find the exact solution. x^2+35=5x

OpenStudy (anonymous):

\[x = \frac{5 \pm \sqrt{25 - 4 \times35} }{2}\]

OpenStudy (anonymous):

\[ x = \frac{ 5\pm i \sqrt{115}}{2} \]

OpenStudy (anonymous):

-5 guys...

OpenStudy (anonymous):

ZhangA: Ishaan is correct \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] a = 1 b = -5 c = 35 \[x = \frac{-(-5) \pm \sqrt{(-5)^2 - 4(1)(35)}}{2(1)} = \frac{5 \pm \sqrt{25 - 140}}{2} = \frac{5\pm \sqrt{-1}\sqrt{115}}{2} = \frac{5\pm i\sqrt{115}}{2} \]

OpenStudy (anonymous):

Sorry, my bad.